5-gram sample of sodium chloride (NaCl) was diluted in 70 grams of water. Determine the percentage composition by mass of the sodium chloride solution.
1
Expert's answer
2019-08-08T04:43:13-0400

"Percentage \space composition = \frac {m \space of \space salt}{m \space of \space solution} \cdot 100 \% = \frac {5g}{5g +70 g} \cdot 100 \% = 6.67 %" %